# FusionPortableV2: A Unified Multi-Sensor Dataset for Generalized SLAM Across Diverse Platforms and Scalable Environments

### v1

@@ 3.1 Suite Setup and Synchronization
Figure 3: Layouts of the platform-specific sensor setup, including different coordinate systems and their relative translation.
3.1.1 Synchronization:
The synchronization process is illustrated in Fig. 2. The FPGA board synchronizes with the pulse-per-second (PPS) signal from its internal GNSS module, producing frequency trigger signals for the IMU, stereo frame cameras, and LiDAR clock alignment. In GPS-denied environments, it utilizes its internal low drift oscillator for synchronization, achieving a time accuracy below $1ms$ for multisensor data collection. For better integration of LiDAR and camera data, we directly synchronize the LiDAR’s phase with the camera’s capture timing. The event cameras, having their own synchronization mechanism, are connected through a designated port. The left event camera, assigned as the master, sends trigger signals to the right camera.
3.1.2 Inertial Measurement Unit:
The STIM300300300300 IMU, a tactical-grade sensor, forms the primary inertial sensor of our system, mounted beneath the LiDAR. Its bias instability for the gyroscope and accelerometer is $0.3^{\circ}/h$ and $0.04mg$, respectively. outputs angular velocity and acceleration measurements at $200Hz$. Other components, including the LiDAR, event cameras, and the $3$DM-GQ7777 INS, also integrate IMUs. Further details are provided in subsequent sections.
@@ 3.1.2 Inertial Measurement Unit:
Our LiDAR choose the OS1-$128$ Gen5555 LiDAR that operates at $10Hz$. It features a built-in IMU capturing gyroscope, acceleration, and magnetometer data and generates four types of images to facilitate the usage of image-based algorithms: range, near-ir, reflectivity, and signal image. Each image measures different properties of the surroundings: (1) range images display the distance of the point from the sensor origin, calculated using the time of flight of the laser pulse; (2) near-ir images capture the strength of sunlight at the $865nm$ light wavelength collected, also expressed in the number of photons detected that was not produced by the sensor’s laser pulse; (3) reflectivity images display the reflectivity of the surface or object that was detected by the sensor; and (4) signal images show the strength of the light returned from the given point, which is influenced by various factors including distance, atmospheric conditions, and objects’ reflectivity.
@@ 3.1.4 Stereo Frame Cameras:
Our setup includes two FLIR BFS-U3333-$31$S4444C global-shutter color cameras for stereo imaging, synchronizing to output images at $1024\times 768$ pixels and $20Hz$. The exposure time $\tau$ for both cameras is manually set as a fixed value. Image timestamps are adjusted by subtracting $0.5\tau$ to approximate the true trigger time. To prevent abrupt changes in color space, the white balance settings are also fixed. Additionally, metadata like exposure time and gain are included for detailed image analysis.
@@ 3.1.5 Stereo Event Cameras:
The system includes two event cameras known for their high temporal resolution, extensive dynamic range, and energy efficiency. With a $346\times 260$ pixel resolution, these cameras come with an internal IMU for high-rate data output. Frame images cannot be synchronized, resulting in $10$-$20ms$ delay. Infrared filters are used to lessen LiDAR light interference. Exposure times are set fixedly, whereas outdoor settings use auto-exposure to maintain image quality under varying light conditions.
3.2 Platform-Specific Sensor Setup
3.2.1 Legged Robot:
We have selected the Unitree A1 quadruped robot as our legged platform, as shown in Fig. 3(c) This robot is equipped with $12$ joint motor encoders and $4$ contact sensors per leg, located at the hip, thigh, calf, and foot. These sensors provide kinematic measurements at a rate of $50Hz$. The MSS is affixed to the robot’s dorsal side and communicates with the kinematic sensors via Ethernet. In addition to the raw sensor measurements, we record metadata for each motor, which includes torque, velocity, position, and temperature, along with kinematic-inertial odometry data.
3.2.2 Unmanned Ground Vehicle:
The MSS is integrated into a four-wheeled Ackerman UGV (see Fig. 3(a), 3(b)), originally designed for logistics transportation (Liu et al., 2021). To optimize signal reception, the dual GNSS antennas of the $3$DM-GQ7777 are positioned at UGV’s rear side. Kinematic data for the UGV is acquired through two incremental rotary encoders, strategically positioned at the center of the rear wheel. These encoders, featuring $1000$ pulses per revolution, produce measurement data at a rate of approximately $100Hz$, which is then recorded.
3.2.3 Vehicle:
As depicted in Fig. 3(d), we follow the KITTI setup (Geiger et al., 2013) by extending the baseline of both the stereo cameras and the dual antenna, with the stereo frame camera having a baseline of $83cm$ and the event camera having a baseline of $73cm$. This extended baseline enhances the accuracy of depth estimation for distant objects, as compared with that in the UGV. The MSS is securely mounted on the vehicle’s luggage rack using a custom-designed aluminum frame.
@@ 3.3.1 Dense RGB Point Cloud Map:
For creating dense point cloud maps of outdoor scenarios, the Leica RTC360 laser scanner was selected, known for its high scanning rate of up to $2$ million points per second and accuracy under $5.3mm$ within a $40m$ radius. Some indoor areas were scanned with the Leica BLK360360360360, which operates at a rate of $0.68$ million points per second and achieves an accuracy of $4mm$ within a $10m$ range. All scans are registered and merged by the Leica Cyclone software, resulting in a dense and precise RGB point cloud map (Fig. 12). This map with the resolution as $8cm$, covering all data collection areas, can be used to evaluate the mapping results of algorithms [^1] https://github.com/JokerJohn/Cloud_Map_Evaluation ranging from model-based (Lin and Zhang, 2022) and learning-based methods (Pan et al., 2024).
@@
For indoor and small-scale outdoor environments, the Leica MS60606060 total station was utilized to measure the GT trajectory of the robot at the $3$-DoF position. As shown in Fig. 4 (a), the tracking prism is placed atop the LiDAR. The GT trajectory was captured at a frequency between $5$-$8Hz$, achieving an accuracy of $1mm$. However, due to occasional instability in the measurement rate, the GT trajectory is resampled at $20Hz$ using linear interpolation for a more consistent evaluation. For further details on this process, please refer to Section 6.2.1.
@@
While the stationary Leica MS60606060 provides accurate measurements, it cannot track the prism when it is occluded or outside the visible range. Consequently, for capturing $6$-DoF GT trajectories in large-scale and outdoor environments with available GNSS satellites, we employ the $3$DM-GQ7777 navigation sensor. This sensor integrates data from its internal dual-antenna RTK-GNSS, which provides raw data at a frequency of $2Hz$, and an IMU, to deliver estimated poses with an output rate of up to $30Hz$. Once the GNSS has initialized with a sufficient satellite lock and RTK is in fixed status, it achieves a positioning accuracy of up to $1.4cm$.
4 Sensor Calibration
We meticulously calibrate the intrinsics of each sensor, their extrinsics, and the time offsets between certain sensors beforehand. The STIM300300300300 IMU’s coordinate system is designated as the body frame, serving as the primary reference for most extrinsic calibrations. For indirectly calibrated sensors, conversion is achieved through matrix multiplication: $\bm{T}^{A}{C}=\bm{T}^{A}{B}\bm{T}^{B}_{C}$. The positioning and orientation of sensors across devices and platforms are illustrated in Fig. 1 and Fig. 3. A summary of the calibration process is provided in Table 3, with detailed results available on our dataset’s website due to page constraints.
4.1 Intrinsic Calibration
We calibrate IMUs and cameras using the off-the-shelf Kalibr toolbox Furgale et al. (2013); Rehder et al. (2016). For wheel encoder intrinsics, such as wheel radius and axle track, we implement the motion-based calibration algorithm outlined in (Jeong et al., 2019). This involves manually maneuver the UGV through significant transformations, as depicted in Fig. 6. We calculate the UGV’s planar motion for each interval $\tau\in[t_{k},t_{k+1}]$ using encoder data to determine linear ($v=(\omega_{l}r_{l}+\omega_{r}r_{r})/2$) and angular ($\omega=(\omega_{l}r_{l}-\omega_{r}r_{r})/b$) velocities. Concurrently, the $3$DM-GQ7777 captures more accurate motion estimates. Intrinsics are then optimized by minimizing the trajectory alignment error between these two trajectories.
@@ 4.1 Intrinsic Calibration
Intrinsics IMU Noisy Density, Random Walk Allen variance analysis Toolbox [^2] https://github.com/ori-drs/allan_variance_ros
@@
Camera Focal Length, Center Point, Distortion Minimize reprojection error (Zhang, 2000) [^3] https://www.mathworks.com/help/vision/camera-calibration.html
Extrinsics IMU-IMU Rotation, Translation Optimization (Rehder et al., 2016) [^4] https://github.com/ethz-asl/kalibr/wiki/Multi-IMU-and-IMU-intrinsic-calibration
IMU-Camera Rotation, Translation, Cons. Time Offset Optimization (Furgale et al., 2013) [^5] https://github.com/ethz-asl/kalibr/wiki/camera-imu-calibration
@@
IMU-Legged Sensors Rotation, Translation Obtained from the CAD Model
Camera-Camera Rotation, Translation Minimize reprojection errors (Zhang, 2000) [^6] https://www.mathworks.com/help/vision/camera-calibration.html
Camera-LiDAR Rotation, Translation Minimize point-to-line and point-to-plane errors (Jiao et al., 2023) [^7] https://github.com/HKUSTGZ-IADC/LCECalib
##\## detailed in the paper
@@ 4.2 Extrinsic Calibration
Extrinsic calibrations, encompassing $6$-DoF transformations and time offsets for IMU-IMU, IMU-camera, IMU-prism, camera-camera, and camera-LiDAR pairs, are typically obtained with off-the-shelf toolboxes. We specifically describe the calibration between the IMU and the prism that defines the reference frame of GT measurements relative to the Leica MS60606060. We design the indirect calibration method since the Leica MS60606060 provides only $3$-DoF and low-rate trajectories. We observed that the prism is visible to infrared cameras in the motion capture room. We place and adjust three tracking markers around the prism to approximate its center, as shown in Fig. 5. We move the handheld device to perform the “$8$”-shape trajectory. Both the motion capture system and LiDAR-inertial odometry (Xu et al., 2022) can estimate trajectories of the prism and STIM300300300300 IMU, respectively. With these trajectories, the IMU-Prism extrinsics are estimated using the algorithm proposed in (Furrer et al., 2018).
Figure 6: Comparision of trajectories: estimated motion by the $3$DM-GQ7777 (red), integration of encoders’ measurement before calibration (green) and after calibration (blue), respectively.
@@ 5.1 Analysis of Platforms Characteristics
(a) Motion Analysis of the Sequence: Handheld_room00
(b) Motion Analysis of the Sequence: Legged_grass00
(c) Motion Analysis of the Sequence: Ugv_parking03
(d) Motion Analysis of the Sequence: Vehicle_highway00
Figure 7: Motion analysis with four mobile platforms in terms of acceleration $[m/s^{2}]$ and angular velocity $[rad/s]$. The left two columns of each row illustrate time-domain data, revealing immediate dynamic behaviors, while the right two columns display frequency-domain data, highlighting predominant motion features. This comparison reveals the robust requirements of general SLAM algorithms on multiple platforms.
@@ 5.1.1 Handheld:
Since the handheld MSS is commonly held by a user, it offers great flexibility for data collection scenarios. The handheld multi-sensor device provides adaptable data collection across diverse settings, akin to market counterparts like the Leica BLK2222GO mobile scanning device, which excels in precision scanning and motion estimates. Therefore, we collect data in scenarios including a laboratory with furniture and dynamic elements, uneven grasslands, an escalator for vertical transitions, and a underground parking lot resembling long tunnels. The device performs motion influenced by the user’s walking or running, sometimes leading to camera shake and rapid directional shifts. Each sequence contains a least one loop. The average movement speed is around $2m/s$.
@@ 5.1.3 Unmanned Ground Vehicle:
The UGV is typically designed for last-mile delivery and navigates middle-scale areas like campuses and factories. Constrained by Ackermann steering geometry, the UGV executes planar and smooth movements in response to the operator’s inputs. Data collection is conducted in various environments, including an outdoor parking lot (lacking structure), a campus, and the challenging transition zones between indoor and outdoor environments (where GNSS signals are unstable). To mimic real-world complexities, commands for sudden stops and $45^{\circ}$ turns are occasionally issued. The UGV can move at speeds of approximately $5m/s$.

### v2

@@ 3.1.1 Synchronization:
Figure 3: Layouts of the platform-specific sensor setup, including different coordinate systems and their relative translation. More detailed and accurate dimensional data are provided in our calibration files.
@@
Our LiDAR choose the OS1-$128$ Gen5555 LiDAR that operates at $10Hz$. It features a built-in IMU capturing gyroscope, acceleration, and magnetometer data and generates four types of images to facilitate the usage of image-based algorithms: range, near-ir, reflectivity, and signal image. Each image measures different properties of the surroundings: (1) range images display the distance of the point from the sensor origin, calculated using the time of flight of the laser pulse; (2) near-ir images capture the strength of sunlight at the $865nm$ light wavelength collected, also expressed in the number of photons detected that were not produced by the sensor’s laser pulse; (3) reflectivity images display the reflectivity of the surface or object that was detected by the sensor; and (4) signal images show the strength of the light returned from the given point, which is influenced by various factors including distance, atmospheric conditions, and objects’ reflectivity. The timestamp of the LiDAR data represents the end of the scanning period.
@@ 3.1.3 Stereo Frame Cameras:
Our setup includes two FLIR BFS-U3333-$31$S4444C global-shutter color cameras for stereo imaging, synchronizing to output images at $1024\times 768$ pixels and $20Hz$. The exposure time $\tau$ for both cameras is manually set as a fixed value. Image timestamps are adjusted by subtracting $0.5\tau$ to properly align image features with IMU measurements since an image is the result of integrating light over the exposure interval (Furgale et al., 2013; Rehder et al., 2016b). To prevent abrupt changes in color space, the white balance settings are also fixed. Additionally, metadata like exposure time and gain are included for image analysis.
@@ 3.1.4 Stereo Event Cameras:
Two event cameras, which are known for their high temporal resolution, extensive dynamic range, and energy efficiency, are used for data collection. These cameras output events data, $346\times 260$ frame images, and high-rate IMU measurements. Frame images cannot be synchronized, resulting in a $10$-$20ms$ delay. Infrared filters are used to lessen LiDAR light interference. Exposure times are set fixedly, whereas outdoor settings use auto-exposure to maintain image quality under varying light conditions.
@@
3.1.5 Inertial Measurement Unit:
The STIM300300300300 IMU, a tactical-grade sensor22endnote: 2https://www.sensonor.com/media/1132/ts1524r9-datasheet-stim300.pdf, serves as the primary inertial sensor of our system, mounted beneath the LiDAR. It has a bias instability of $0.3^{\circ}/h$ for the gyroscope and $0.04mg$ for the accelerometer. The sensor outputs angular velocity and acceleration measurements at $200Hz$. Other components, including the LiDAR, event cameras, and the $3$DM-GQ7777 Inertial Navigation System (INS), are also integrated with IMUs. Further details are provided in subsequent sections.
3.2 Platform-Specific Sensor Setup
3.2.1 Legged Robot:
We have selected the Unitree A1 quadruped robot as our legged platform, as shown in Fig. 3(c). This robot is equipped with $12$ joint motor encoders and $4$ contact sensors per leg, located at the hip, thigh, calf, and foot. These sensors provide kinematic measurements at a rate of $50Hz$. The MSS is affixed to the robot’s dorsal side and communicates with the kinematic sensors via Ethernet. In addition to the raw sensor measurements, we record metadata for each motor, which includes torque, velocity, position, and temperature, along with kinematic-inertial odometry data.
3.2.2 Unmanned Ground Vehicle:
The MSS is integrated into a four-wheeled Ackerman UGV (see Fig. 3(a), 3(b)), originally designed for logistics transportation (Liu et al., 2021). To optimize signal reception, the dual GNSS antennas of the INS are positioned at UGV’s rear side. Kinematic data for the UGV is acquired through two incremental rotary encoders, strategically positioned at the center of the rear wheel. These encoders, featuring $1000$ pulses per revolution, produce measurement data at a rate of approximately $100Hz$, which is then recorded.
3.2.3 Vehicle:
As depicted in Fig. 3(d), we follow the KITTI setup (Geiger et al., 2013) by extending the baseline of both the stereo cameras and the dual antenna, with the stereo frame camera having a baseline of $83cm$ and the event camera having a baseline of $73cm$. This extended baseline enhances the accuracy of depth estimation for distant objects, as compared with that in the UGV. The MSS is mounted on the vehicle’s luggage rack using a custom-designed aluminum frame.
@@ 3.3.1 Dense RGB Point Cloud Map:
For creating dense point cloud maps of outdoor scenarios, the Leica RTC360 laser scanner was selected, because of its high scanning rate of up to $2$ million points per second and accuracy under $5.3mm$ within a $40m$ radius. Some indoor areas were scanned with the Leica BLK360360360360, which operates at a rate of $0.68$ million points per second and achieves an accuracy of $4mm$ within a $10m$ range. All scans are registered and merged by the Leica Cyclone software33endnote: 3https://leica-geosystems.com/products/laser-scanners/software/leica-cyclone/leica-cyclone-register-360, resulting in a dense and precise RGB point cloud map (Fig. 12). This map with the resolution as $8cm$, covering all data collection areas, can be used to evaluate the mapping results of algorithms44endnote: 4https://github.com/JokerJohn/Cloud_Map_Evaluation ranging from model-based (Lin and Zhang, 2022) and learning-based methods (Pan et al., 2024).
@@
For indoor and small-scale outdoor environments, the Leica MS60606060 total station55endnote: 5https://leica-geosystems.com/en-us/products/total-stations/multistation/leica-nova-ms60 was utilized to measure the GT trajectory of the robot at the $3$-DoF position. As shown in Fig. 4 (a), the tracking prism is placed atop the LiDAR. The GT trajectory was captured at a frequency between $5$-$8Hz$, achieving an accuracy of $1mm$. However, due to occasional instability in the measurement rate, the GT trajectory is resampled at $20Hz$ using cubic spline interpolation for a more consistent evaluation. For further details on this process, please refer to Section 6.2.1.
@@
While the stationary Leica MS60606060 provides accurate measurements, it cannot track the prism when it is occluded or outside the visible range. Consequently, we employ the INS to capture $6$-DoF GT trajectories in large-scale and outdoor environments with available GNSS satellites. This sensor integrates data from its internal dual-antenna RTK-GNSS, which provides raw data at a frequency of $2Hz$, and an IMU, to deliver estimated poses with an output rate of up to $30Hz$. Before commencing data collection, we ensure that the GNSS has initialized with a sufficient satellite lock and the RTK is in a fixed status, typically achieving a positioning accuracy of up to $1.4cm$. This initialization process usually takes $1$-$3min$ in outdoor. To maintain the reliability of our ground truth data, we adhere to strict criteria. Firstly, the INS filter must be in a stable navigation status. Secondly, both dual-antenna GNSS must be in RTK-fixed status, receiving signals from at least $20$ satellites. Lastly, the positional error covariance must have converged to an optimal state.
@@
Intrinsics IMU Noisy Density, Random Walk Allen variance analysis Toolbox66endnote: 6https://github.com/ori-drs/allan_variance_ros
@@
Camera Focal Length, Center Point, Distortion Minimize reprojection error (Zhang, 2000)77endnote: 7https://www.mathworks.com/help/vision/camera-calibration.html
Extrinsics IMU-IMU Rotation, Translation Optimization (Rehder et al., 2016a)88endnote: 8https://github.com/ethz-asl/kalibr/wiki/Multi-IMU-and-IMU-intrinsic-calibration
IMU-Camera Rotation, Translation, Cons. Time Offset Optimization (Furgale et al., 2013)99endnote: 9https://github.com/ethz-asl/kalibr/wiki/camera-imu-calibration
@@
IMU-Legged Sensors Rotation, Translation Obtained from the CAD Model 1010endnote: 10https://github.com/unitreerobotics/unitree_ros/blob/master/robots/a1_description/urdf/a1.urdf
Camera-Camera Rotation, Translation Minimize reprojection errors (Zhang, 2000)
Camera-LiDAR Rotation, Translation Minimize point-to-line and point-to-plane errors (Jiao et al., 2023)$1$endnote: 11https://github.com/HKUSTGZ-IADC/LCECalib
##\## detailed in Section 4.2 of the paper
4 Sensor Calibration
@@
We meticulously calibrate the intrinsics of each sensor, their extrinsics, and the time offsets between certain sensors beforehand. The STIM300300300300 IMU’s coordinate system is designated as the body frame, serving as the primary reference for most extrinsic calibrations. For indirectly calibrated sensors, conversion is achieved through matrix multiplication: $\mathbf{T}^{A}_{C}=\mathbf{T}^{A}_{B}\mathbf{T}^{B}_{C}$. The positioning and orientation of sensors across devices and platforms are illustrated in Fig. 1 and Fig. 3. A summary of the calibration process is provided in Table 3. We show detailed results and guidelines for good calibration on the dataset’s website due to page constraints.
4.1 Intrinsic Calibration
We calibrate IMUs and cameras using the off-the-shelf Kalibr toolbox (Furgale et al., 2013; Rehder et al., 2016a). For wheel encoder intrinsics, such as wheel radius and axle track, we implement the motion-based calibration algorithm outlined in (Jeong et al., 2019). This involves manually maneuver the UGV through significant transformations, as depicted in Fig. 6. We calculate the UGV’s planar motion for each interval $\tau\in[t_{k},t_{k+1}]$ using encoder data to determine linear $v=(\omega_{l}r_{l}+\omega_{r}r_{r})/2$ and angular $\omega=(\omega_{l}r_{l}-\omega_{r}r_{r})/b$ velocities. Concurrently, the INS captures more accurate motion estimates, and intrinsics are then optimized by minimizing the trajectory alignment error between these two trajectories.
@@ 4.2 Extrinsic Calibration
Extrinsic calibrations, encompassing $6$-DoF transformations and time offsets for IMU-IMU, IMU-camera, IMU-prism, camera-camera, and camera-LiDAR pairs, are typically obtained with off-the-shelf toolboxes. We specifically describe the calibration between the IMU and the prism that defines the reference frame of GT measurements relative to the total station (Leica MS60606060). We design the indirect calibration method since the total station provides only $3$-DoF and low-rate trajectories. We observed that the prism is visible to infrared cameras in the motion capture room. We place and adjust three tracking markers around the prism to approximate its center, as shown in Fig. 5. We move the handheld device to perform the “$8$”-shape trajectory. Both the motion capture system and LiDAR-inertial odometry (Xu et al., 2022) can estimate trajectories of the prism and STIM300300300300 IMU (body frame), respectively. With these trajectories, the IMU-Prism extrinsics are estimated using the hand-eye calibration algorithm (Furrer et al., 2018).
Figure 6: Comparison of trajectories: estimated motion by the INS ($3$DM-GQ7777) (red), integration of encoders’ measurements before calibration (green), and after calibration (blue) using the sequence Ugv_parking00 for calibration. The ATE of the trajectory alignment is $0.98m$.
@@ 5.1 Analysis of Platforms Characteristics
(a) Motion Analysis of the Sequence: Handheld_room00.
(b) Motion Analysis of the Sequence: Legged_grass00.
(c) Motion Analysis of the Sequence: Ugv_parking00.
(d) Motion Analysis of the Sequence: Vehicle_highway00.
(e) Linear Velocity of Above Sequences: Handheld_room00, Legged_grass00, Ugv_parking00, Vehicle_highway00.
Figure 7: Motion analysis with four mobile platforms in terms of linear acceleration $[m/s^{2}]$, angular velocity $[rad/s]$ and velocity $[m/s]$. We use measurements from STIM-$300$ to get linear acceleration and angular velocity and SLAM results to get rough velocity. (a)-(d): The left two columns of each row illustrate time-domain data, revealing immediate dynamic behaviors, while the right two columns display frequency-domain data, highlighting predominant motion features. Each line in the legend represents linear acceleration and angular velocity of each axis: $X$- (Red), $Y$- (Green), and $Z$- (Purple). (e) The linear velocity of these sequences.
@@ 5.1.1 Handheld:
Since the handheld MSS is commonly held by a user, it offers flexibility for data collection scenarios. The handheld multi-sensor device provides adaptable data collection across diverse settings, akin to market counterparts like the Leica BLK2222GO mobile scanning device, which excels in precision scanning and motion estimates. Therefore, we collect data in scenarios including a laboratory with furniture and dynamic elements, uneven grasslands, an escalator for vertical transitions, and an underground parking lot resembling long tunnels. The device performs motion influenced by the user’s walking or running, sometimes leading to camera shake and rapid directional shifts. Each sequence contains at least one loop. The average movement speed is around $2m/s$.
@@ 5.1.3 Unmanned Ground Vehicle:
The UGV is typically designed for last-mile delivery and navigates middle-scale areas like campuses and factories. Constrained by Ackermann steering geometry, the UGV executes planar and smooth movements in response to the operator’s inputs. Data collection is conducted in various environments, including an outdoor parking lot (open space), a campus, and the challenging transition zones between indoor and outdoor environments (where GNSS signals are unstable). To mimic real-world complexities, commands for sudden stops and $45^{\circ}$ turns are occasionally issued. The UGV can move at speeds of approximately $5m/s$.